A Complete Bibliography of ACM Transactions on Modeling and Performance Evaluation of Computing

Systems (TOMPECS)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

02 March 2022 Version 1.05

Title word cross-reference

3 [YLL20]. d [HV19, VV19]. N [WZK^{+}19].

-Choices [VV19]. -Tier [WZK^{+}19].

1 [GHV20].

802.15.4 [AM17]. 8th [KS18].

Accelerators [GLL^{+}21]. Access [NCF^{+}17]. Access-Time-Aware [NCF^{+}17]. Accuracy [MIS21]. ACM [KS18]. ACM/SPEC [KS18]. across [LKC^{+}21]. Action [KHN^{+}18]. Adaptive [MSN^{+}21].

Admission [RAMB20]. Ads [VGCL20]. Advance [SS17]. Advertisements [VGCL20]. Affecting [WLU18]. Algorithm [VV19]. Algorithms [NCF^{+}17, QE21]. Allocating [WDGC19]. Allocation [MRS20, RAMB20, SSB^{+}20, TMASA16, WZK^{+}19]. Allocations [FPW17]. Amazon [WLU18]. AMIR [KKR19]. Analysis [CCH^{+}16, DGRL16, DWS17, EJ21, FGK^{+}21, FGRT16, GLM16, LRS18, LLW^{+}19, PFK18, PPP^{+}17, PTA^{+}20, RBL20, SKV21, WLU18, XLT16]. Analytic [AM20, KKR19, YMRS16]. Answer [KSM^{+}17]. App [PPP^{+}17]. Application [IADB19, PTA^{+}20, SSB^{+}20, WZK^{+}19]. Application- [SSB^{+}20]. Applications [DWS17, DD17, DD18, FA19, GHV20, MRH18, PAEÅ^{+}16, ZWHD16]. Apportionment [VNTA16]. Approach [GLM16, SSM20]. Approximation [JNT18]. Approximations [IAV16]. AQM [DGRL16]. Architecture [MRH18]. Architectures [MRH18]. ARM [AFGR18]. ARM-based [AFGR18]. Arrivals [WXL^{+}19]. Attacks [GG21]. Attribution [NWK^{+}16]. Auto [NXL17, PAEÅ^{+}16]. Auto-Scaling [NXL17, PAEÅ^{+}16]. Automatic [AM20]. Autoscalers [IAEH^{+}18]. Aware [CZCC19, CDPN21, MC16, NCF^{+}17, SSB^{+}20].

Balancing [YYX^{+}19]. Bargaining [LLTL18]. Based [AM17, DD17, FGK^{+}21, LLTL18, NWK^{+}16, NASD21, RAMB20, VV19, YLN^{+}17, AFGR18]. Batched [GLL^{+}21]. Beacon [AM17]. Beacon-Enabled [AM17]. Behavioral [AM17]. Benefits [LB16]. Big [MRH18]. Block [LSC^{+}20]. Bloom [FB16]. Bounded [SSM20]. Broadcast [SLH19]. Bundle [QE21]. Burstiness [KKR19]. Buses [CCH^{+}16]. Buying [YLN^{+}17].

Cache [FA19, JNT18, NCF^{+}17]. Caches [LB16]. Caching [DMD^{+}21, GLM16, QE21]. Calls [CCY^{+}18]. Capacity [FPW17, LS17]. Case [NT16]. Center [CZCC19, EJ21]. Centers [FGRT16, ZRWL16]. Centric [CCY^{+}18]. Chains [EJ21]. Challenges [HBK^{+}18]. Channel [RAMB20]. Characteristics [BCG19, YLL20, ZWHD16]. Characterization [MRH18, NASD21]. Chief [Gol21]. Chip [LS17]. Choices [VV19, YYX^{+}19]. Chunked [SLH19]. Churn [WXL^{+}19]. Class [AFGR18, GHV20]. Cloud [AAA21, AAL^{+}17, CPFC20, DD17, GSS16, HBK^{+}18, JSW17, LZL^{+}19, LLTL18, MC16, NWK^{+}16, PAEÅ^{+}16, WZK^{+}19, WDGC19, YLN^{+}17, ZLW18]. Cloud-Based [DD17]. CloudHeat [CZL^{+}18]. Clouds [JLZ20]. Clustering [LLW^{+}19, VS19]. Cocoa [YLN^{+}17]. Coded [GG21]. Codes [CCH^{+}16]. Coding [CCH^{+}16, LRS18, SLH19]. Collection [VV19]. Colocation [ZRWL16]. Colored [AM17]. Combinatorial [EJ21]. Combined [LS17]. Community [RSS18]. Comparative [LZL^{+}19]. Competing [NBP19]. Completeness [GHV20]. Complex [IAEH^{+}18]. Compression [FB16]. Computing [DD17, FGRT16, LLTL18, SKV21, WJW19, WDGC19, YLN^{+}17, ZLW18]. Concurrency [LS17]. Conference [KS18]. Configuration [BJLM16]. Congestion [DGRL16]. Considerations [VNTA16]. Consistent [WXL^{+}19]. Contact [BBPC17]. Container [YLN^{+}17]. Container-Based [YLN^{+}17]. Content [GM19, NXL17, PFK18]. Contention [MC16]. Contention-Aware [MC16]. Control [CCY^{+}18, DGRL16, RAMB20]. Control-Theoretic [DGRL16]. Controlling [FPW17]. Convergence [JNT18]. Copula [DWS17]. Core [LS17, LCD^{+}17, MRH18, NSMA19]. Correlated [NT16]. Cost [WLU18]. Cost-Effective [WLU18]. Costs [NWK^{+}16]. Covert [JNT21]. Crowd [NTLM18]. Crowdsourcing [XLT16]. CTMCs [SSM20]. CTMDPs [SSM20]. Cycle [JNT21]. Cycling [BBPC17].

D [YLL20]. DAG [FGK^{+}21]. DAG-Based [FGK^{+}21]. Darknets [RS16]. Data [CZCC19, CDPN21, EJ21, FGRT16, IADB19, KSM^{+}17, LRS18, LKC^{+}21, MSN^{+}21, MRH18, WCKN18, ZRWL16]. Data-Intensive [KSM^{+}17, MSN^{+}21]. Databases [MC16]. Datacenter [CZL^{+}18]. Dead [RS16]. Dealing [RS16]. Decisions [RAMB20]. Deferrals [FPW17]. Deficit [DD18]. Delay [SLH19]. Demand [GM19, WDGC19, ZRWL16, ZLW18]. Demands [WCKN18]. Dependability [HBK^{+}18]. Dependence [DWS17]. Derivation [AM20]. Design [HBK^{+}18, LS17, LKC^{+}21, XLT16]. Detecting [KSSO16]. Detection [PPIR19]. discharge [RBL20]. Discrete [GLL^{+}21]. Discrete-Time [GLL^{+}21]. Discriminatory [IAV16]. Disk [CSS^{+}18]. Dissemination [NT16]. Distributed [FGK^{+}21, GG21, LSC^{+}20, QE21, WXL^{+}19, ZRWL16]. Distributions [AP16]. Diverse [LKC^{+}21]. DTNs [NBP19]. Duty [BBPC17]. Dynamic [DD18, GHV20, MRS20, NT16, RAMB20, TMASA16, YLN^{+}17]. Dynamics [LZW^{+}16].

EC2 [WLU18]. Economics [AAL^{+}17]. Ecosystem [PPP^{+}17]. Edge [NMK21]. Editor [Gol21]. Editor-in-Chief [Gol21]. Effect [GM19, LS17]. Effective [WLU18]. Effectiveness [CPFC20]. Efficiency [DD18, LZL^{+}19]. Efficient [CZL^{+}18, JSW17, LSC^{+}20, LKC^{+}21, RS16, WJW19, YYX^{+}19]. Embedded [NSMA19]. Emergency [ZRWL16, ZLW18]. Emerging [HBK^{+}18]. Empirical [WLU18]. Employing [LB16]. Enabled [AM17]. End [SSB^{+}20]. End-Host [SSB^{+}20]. Ends [RS16]. Endurance [VV19]. Energy [HPK16, LCD^{+}17, RBL20, VNTA16]. EnergyQARE [CZCC19]. Engagement [PTA^{+}20]. Engineering [KS18]. Ensuring [PFK18]. Enterprise [SSB^{+}20]. Environment [MSN^{+}21]. Environments [MC16]. EP [HPK16]. EQ [CCY^{+}18]. Estimating [BJLM16]. Estimation [AP16, VS19]. Evaluating [LS17, PPIR19]. Evaluation [IAEH^{+}18, KKRK19, LKC^{+}21, MH20, PAEÅ^{+}16, TMASA16]. Expansion [YYX^{+}19]. Experiences [KHN^{+}18]. Experimental [IAEH^{+}18]. Experiments [LLW^{+}19]. Exploit [GLL^{+}21]. Exploring [YLL20].

Fair [NWK^{+}16]. Fairness [VNTA16]. False [FB16]. False-Positive [FB16]. Fast [LCD^{+}17]. Features [WLU18]. Field [LRS18, LXG^{+}18]. FIFO [JNT21]. File [NT16, QE21, SLH19]. File-Bundle [QE21]. File-Transfer [SLH19]. Filters [FB16]. First [DD18]. Fixed [NBP19]. Flash [VV19]. Flash-Based [VV19]. Flow [AP16, Var18]. Fog [FGRT16]. Forbidden [CCH^{+}16]. Fork [MRS20]. Framework [FGRT16, LXW^{+}17, PAEÅ^{+}16, PPIR19, WDGC19, YMRS16]. Function [EJ21]. Future [LCD^{+}17].

G [GHV20]. Game [LLTL18]. Game-Based [LLTL18]. Games [LXG^{+}18, SS17]. Garbage [VV19]. Generalization [FA19, QE21]. Generic [EJ21]. Geo [ZRWL16]. Geo-Distributed [ZRWL16]. Geographic [Var18]. Go [PTA^{+}20]. GPSonflow [Var18]. Graph [RSS18]. Graphs [NT16]. Grid [CZCC19]. Group [YLN^{+}17]. Guarantees [LLTL18].

Harvesting [CZL^{+}18]. HDD [YMRS16]. Heat [CZL^{+}18]. High [CCH^{+}16]. High-Speed [CCH^{+}16]. hop [NBP19]. Horizontal [JLZ20]. Host [NASD21, SSB^{+}20]. Host-Based [NASD21]. Hosting [NMK21]. HPC [CPFC20]. HTTP [CSS^{+}18]. Hypervisor [NASD21]. Hysteresis [DMD^{+}21].

ICPE [KS18]. Identical [HV19]. Identical/Independent [HV19]. Identification [GG21]. IEEE [AM17]. iModel [AM20]. Impact [PTA^{+}20]. Impacts [BBPC17]. Improving [KKR19]. In-Memory [MC16]. Incentive [NBP19, VNTA16, XLT16, ZRWL16]. Increasing [CSS^{+}18]. Incremental [YYX^{+}19]. Independent [HV19, JNT18, NT16]. Inference [WCKN18]. Insertion [LXW^{+}17]. Instance [WLU18]. Intensive [KSM^{+}17, MSN^{+}21]. Interaction [VGCL20]. Interface [LKC^{+}21]. International [KS18]. Internet [BCG19]. Introduced [LLW^{+}19]. Introduction [MH20, TW16]. Investigating [BCG19]. IOTA [FGK^{+}21]. Issue [KS18]. Item [NTLM18].

Join [MRS20]. Joint [AAA21].

Large [LRS18, SKV21, WJW19, ZWHD16]. Large-Scale [WJW19, ZWHD16, SKV21]. Largest [DD18]. Latency [JSW17]. Ledger [FGK^{+}21]. Level [LZL^{+}19, MRH18]. Limit [JLZ20]. Linear [SLH19]. Links [NT16]. List [Abi18, Abo17, Ano16, BW19]. Little [MRH18]. Load [YYX^{+}19]. Load-Balancing [YYX^{+}19]. Lose [BBPC17]. Low [DGRL16]. Low-Priority [DGRL16]. LRU [FA19, JNT18]. Lyapunov [SSM20].

M [GHV20]. M/G/1 [GHV20]. Machine [NASD21]. Malicious [GG21]. Managed [LB16]. Management [CDPN21, LSC^{+}20, LCD^{+}17, YMRS16]. Managing [HPPQ19, KSM^{+}17]. Mansard [MIS21]. Many [LS17, LCD^{+}17]. Many-Core [LS17, LCD^{+}17]. Mapping [CDPN21, NSMA19]. MapReduce [PPIR19]. Market [CZL^{+}18]. Markov [DWS17]. Mean [LRS18, LXG^{+}18]. Mean-Field [LRS18]. Measurement [PPP^{+}17]. Measurements [KHN^{+}18]. Mechanism [CZL^{+}18, CCY^{+}18, ZRWL16, ZLW18]. Mechanisms [CSS^{+}18, PPIR19, XLT16]. Memory [CDPN21, LS17, MC16]. Message [Gol21]. Method [KKR19]. Methodology [WCKN18]. Metric [HBK^{+}18]. Mining [NASD21]. Mobile [PPP^{+}17]. Mode [AM17]. Model [AM17, MIS21, RBL20].

Modeling [CPFC20, GLL^{+}21, LLW^{+}19, MSN^{+}21, PPP^{+}17, VGCL20]. Models [AM20, BJLM16]. Modulated [DWS17]. Monitoring [IADB19]. Multi [NSMA19]. Multi-Core [NSMA19]. Multithreaded [SKV21].

Net [AM17]. Network [LZL^{+}19, LZW^{+}16, SLH19]. Network-Level [LZL^{+}19]. Networks [BJLM16, BBPC17, CPFC20, EJ21, LXG^{+}18, PFK18, SSB^{+}20]. Neural [CPFC20]. NFV [GLL^{+}21]. Nice [Var18]. Node [GG21]. Non [WXL^{+}19]. Non-Stationary [WXL^{+}19]. Nudge [LXG^{+}18].

Obtaining [KSM^{+}17]. Offloading [FGRT16]. Offs [HPK16]. On-Demand [WDGC19]. Online [CZL^{+}18, CDPN21, KSM^{+}17, NSMA19, QE21, VGCL20, XLT16, ZLW18]. OpenACC [LB16]. OpenFOAM [LXW^{+}17]. Opportunistic [BBPC17, PFK18]. Opportunities [LB16]. Optane [YLL20]. Optimal [QE21, SLH19, Var18]. Optimality [DD18]. Optimization [AAA21, FB16, VGCL20]. Optimizing [WZK^{+}19]. Output [KKRK19]. Overlap [CCH^{+}16].

Pacing [SSB^{+}20]. Packet [LLW^{+}19]. Page [TMASA16]. PageRank [VS19]. Paper [KS18]. Parallel [KKRK19, WJW19, ZWHD16]. Parameterized [GHV20]. Participation [CZCC19]. Paths [BCG19]. Peak [NWK^{+}16]. Peak-Based [NWK^{+}16]. PEAS [PAEÅ^{+}16]. Performance [AM20, FGK^{+}21, GLM16, HPK16, HV19, HBK^{+}18, IAEH^{+}18, IADB19, KKRK19, KS18, LLTL18, MSN^{+}21, NBP19, PAEÅ^{+}16, SKV21, TMASA16, YLL20]. Periodic [WXL^{+}19]. Persistent [PFK18].

Personalized [VS19]. Petri [AM17]. PETSc [LXW^{+}17]. Placement [MC16, VGCL20]. Poisson [DWS17]. Policies [GHV20, VNTA16]. Policy [FA19]. Pollution [GG21]. Positioning [Var18]. Positive [FB16]. Power [CPFC20, KHN^{+}18, LCD^{+}17]. Prefetching [CSS^{+}18]. PREFiguRE [YMRS16]. Pricing [NWK^{+}16]. Prioritization [DD18]. Priority [DGRL16, GHV20]. Probability [BJLM16, FB16]. Process [BBPC17, DWS17]. Processes [JNT18]. Processing [GLL^{+}21]. Processor [IAV16]. Procurement [WLU18]. Production [IADB19]. Provision [CZCC19, RAMB20]. Provisioning [KKRK19, MSN^{+}21]. Public [JLZ20].

QEST [MH20]. QMLE [WCKN18]. QoE [CCY^{+}18]. QoE-Centric [CCY^{+}18]. QoS [CZCC19, PTA^{+}20, RAMB20]. QoS-Aware [CZCC19]. Quality [AAA21, KSM^{+}17, NTLM18]. Quantifying [HBK^{+}18]. Quantitative [LZL^{+}19, MH20]. Queue [HPK16, IAV16]. Queueing [RBL20, WCKN18]. Queues [GHV20, MRS20].

Random [SLH19]. RAPL [KHN^{+}18]. Rate [CCH^{+}16, CCY^{+}18]. Reachability [SSM20]. Real [DD17, DD18]. Real-Time [DD17, DD18]. Recommendations [GM19, KSSO16]. Reducing [KKR19]. Reduction [JSW17]. Redundancy [HV19, JSW17]. Regulation [CZCC19]. Reinforcing [MIS21]. Relays [NBP19]. Reliability [EJ21]. Reliable [NSMA19]. Renting [NMK21]. Replacement [FA19]. Replicas [HV19]. Replication [LRS18, LSC^{+}20, NXL17, WJW19]. Reprioritization [DGRL16]. Reputation [XLT16]. Request [JNT18]. Reservation [SS17]. Reserve [CZCC19]. Resource [MSN^{+}21, MRS20, NXL17, SSB^{+}20, WLU18, WZK^{+}19]. Resources [NMK21]. Response [HPPQ19, ZRWL16, ZLW18]. Responsiveness [KKR19]. Reviewers [Abi18, Abo17, Ano16, BW19]. Reward [NBP19]. Role [DMD^{+}21, VS19]. Roofline [MIS21]. Roofs [MIS21]. Routers [LLW^{+}19]. Routing [RS16, YYX^{+}19].

Sampled [AP16]. Sampling [WXL^{+}19]. Scalability [WZK^{+}19]. Scalable [SSB^{+}20]. Scale [AFGR18, WJW19, ZWHD16, SKV21]. Scale-Out [AFGR18]. Scale-Up [AFGR18]. Scaling [JLZ20, NXL17, PAEÅ^{+}16]. Scheduling [DD17, GSS16, LLTL18, SLH19]. Scheme [FGRT16, NBP19]. Scoring [NTLM18]. Searching [RSS18]. Section [MH20]. Selected [KS18]. Selecting [NTLM18]. Self [LZW^{+}16, RBL20]. Self-discharge [RBL20]. Self-Similarity [LZW^{+}16]. Semi [NSMA19]. Series [NT16]. Server [AFGR18, CSS^{+}18, JNT21, MRH18, WDGC19]. Server-Class [AFGR18]. Service [CZCC19, EJ21, FPW17, NMK21, WCKN18]. Services [KSM^{+}17, LZL^{+}19, WDGC19]. Serving [GM19]. Sharding [HPPQ19]. Sharing [CDPN21, IAV16]. Sharing-Aware [CDPN21]. Should [PTA^{+}20]. Similarity [LZW^{+}16]. Simple [LKC^{+}21]. Simulation [ZWHD16]. Single [JNT21, RSS18]. Smart [CZCC19]. Snooze [BBPC17]. Social [LZW^{+}16]. Societal [LXG^{+}18]. SoCs [AFGR18]. Soft [DD17, DD18, WZK^{+}19]. Software [LB16]. Software-Managed [LB16]. Sojourn [IAV16]. Some [GHV20]. SORT [NSMA19]. Sparse [NXL17]. SPEC [KS18]. Special [KS18, MH20]. Speed [CCH^{+}16]. Sponsored [KSSO16]. Spot [WLU18, WDGC19]. SSDs [TMASA16, VV19]. Stability [PFK18]. Stall [AAA21]. Stationary [JNT18, WXL^{+}19]. Statistical [WCKN18, ZWHD16]. Stay [PTA^{+}20]. Stealing [JNT21, SKV21]. Stochastic [PFK18]. Storage [GG21, LRS18, LZL^{+}19, LSC^{+}20, LKC^{+}21, NXL17, RBL20, Var18, YLL20]. Storms [GSS16]. Straggler [PPIR19, WJW19]. Strategies [PAEÅ^{+}16, TMASA16, YLN^{+}17]. Streaming [AAA21, CSS^{+}18, FA19, IADB19]. Streams [GSS16]. Structure [DWS17]. Structured [FB16]. Study [AFGR18, LZL^{+}19, WZK^{+}19]. Support [DD17]. Synchronization [KKRK19]. System [IADB19, MRH18]. Systems [DMD^{+}21, DD17, GG21, GLM16, JSW17, KKRK19, LRS18, LXG^{+}18, LCD^{+}17, MH20, NSMA19, NXL17, RBL20, WXL^{+}19, XLT16].

Tails [HPPQ19]. Task [NSMA19]. Taxonomy [HBK^{+}18]. Techniques [JSW17]. Technology [YLL20]. Temporal [DWS17, VGCL20]. Tenants [NWK^{+}16]. Theoretic [DGRL16]. Thread [CDPN21]. Throughput [CSS^{+}18]. Tier [WZK^{+}19]. Time [DD17, DD18, GLL^{+}21, HPPQ19, IAV16, NCF^{+}17, SSM20, WDGC19]. Time-Bounded [SSM20]. Top [NTLM18]. Top-Quality [NTLM18]. Trace [NASD21]. Trade [HPK16]. Trade-Offs [HPK16]. Traffic [AP16, YYX^{+}19]. Transfer [LKC^{+}21, SLH19]. Transmission [BJLM16]. Tree [FB16]. Tree-Structured [FB16]. Truthful [ZRWL16]. TTL [JNT18]. Two [GHV20, NBP19]. Two-Class [GHV20]. Two-hop [NBP19].

Unified [GLM16]. Unit [CDPN21]. Unknown [GM19]. User [DD18, PTA^{+}20, SSB^{+}20, VGCL20]. User-aware [SSB^{+}20]. Using [CDPN21, FPW17, KHN^{+}18, NASD21, SSB^{+}20, ZWHD16, SLH19].

Validating [ZWHD16]. Variability [FPW17]. Versus [LRS18]. via [HPK16, PFK18, YYX^{+}19]. VidCloud [AAA21]. Video [AAA21, CSS^{+}18, FA19, NXL17]. Virtual [NASD21]. Virtualized [MSN^{+}21]. VoIP [CCY^{+}18]. vs [AFGR18].

Wireless [BJLM16]. Work [SKV21]. Workflows [IAEH^{+}18]. Workload [MC16, NASD21]. Workloads [AFGR18, MSN^{+}21]. Worth [CPFC20].

Xpoint [YLL20].

YouTube [PTA^{+}20].

References

Al-Abbasi:2021:VJS

[AAA21] Abubakr O. Al-Abbasi and Vaneet Aggarwal. VidCloud: Joint stall and quality optimization for video streaming over cloud. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 5(4):17:1–17:32, March 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                                   tronic). URL https://dl.acm.
                                                   org/doi/10.1145/3442187.

Anselmi:2017:EC

[AAL^{+}17] Jonatha Anselmi, Danilo Ardagna, John C. S. Lui, Adam Wierman, Yunjian Xu, and Zichao Yang. The economics of the cloud. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(4):18:1–18:??, December 2017. CODEN ????

ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3086574.

Abid:2018:LR

[Abi18] Amine Abid. List of reviewers. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(4):21:1–21:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3271430.

Abouzeid:2017:LR

[Abo17] Alhussein Abouzeid. List of reviewers. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(4):23:1–23:??, December 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3162084.

Azimi:2018:SVS

[AFGR18] Reza Azimi, Tyler Fox, Wendy Gonzalez, and Sherief Reda. Scale-out vs scale-up: A study of ARM-based SoCs on serverclass workloads. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3 (4):18:1–18:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/citation.cfm?id=3232162.

Alves:2017:BMI

[AM17] Renan C. A. Alves and Cíntia B. Margi. Behavioral model of IEEE 802.15.4 beacon-enabled mode based on colored Petri net. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(4):20:1–20:??, December 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3115389.

Awad:2020:IAD

[AM20] Mahmoud Awad and Daniel A. Menascé. iModel: Automatic derivation of analytic performance models. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5 (2):7:1–7:30, April 2020. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3374220.

Anonymous:2016:LR

[Ano16] Anonymous. List of reviewers. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 1(4):20:1–20:2, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2989212.

Antunes:2016:EFD

[AP16] Nelson Antunes and Vladas Pipiras. Estimation of flow distributions from sampled traffic. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 1(3):11:1–11:28, May 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2891106.

Biondi:2017:WYL

[BBPC17] Elisabetta Biondi, Chiara Boldrini, Andrea Passarella, and Marco Conti. What you lose when you snooze: How duty cycling impacts on the contact process in opportunistic networks. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(4):22:1–22:??, December 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3149007.

Bakhshaliyev:2019:ICI

[BCG19] Khalid Bakhshaliyev, Muhammed Abdullah Canbaz, and Mehmet Hadi Gunes. Investigating characteristics of Internet paths. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 4(3):16:1–16:??, September 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3342286.

Bermolen:2016:ETP

[BJLM16] Paola Bermolen, Matthieu Jonckheere, Federico Larroca, and Pascal Moyal. Estimating the transmission probability in wireless networks with configuration models. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(2):9:1– 9:23, June 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2858795.

Borst:2019:LR

[BW19] Sem Borst and Carey Williamson. List of reviewers. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4 (4):23:1–23:2, December 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/abs/10.1145/3369841.

Chang:2016:CRA

[CCH^{+}16] Cheng-Shang Chang, Jay Cheng, Tien-Ke Huang, Duan-Shin Lee, and Cheng-Yu Chen. Coding rate analysis of forbidden overlap codes in high-speed buses. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1 (2):8:1–8:25, June 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

                                        http://dl.acm.org/citation.
                                        cfm?id=2846091.

Chu:2018:EQC

[CCY^{+}18] Cing-Yu Chu, Shannon Chen, Yu-Chuan Yen, Su-Ling Yeh, Hao-Hua Chu, and Polly Huang. EQ: A QoE-centric rate control mechanism for VoIP calls. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 3(1):4:1–4:??, February 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3170430.

Cruz:2021:OTD

[CDPN21] Eduardo H. M. Cruz, Matthias Diener, Laércio L. Pilla, and Philippe O. A. Navaux. Online thread and data mapping using a sharing-aware memory management unit. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(4): 16:1–16:28, March 2021. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3433687.

Costa:2020:ENN

[CPFC20] Georges DA Costa, Jean-Marc Pierson, and Leandro Fontoura-Cupertino. Effectiveness of neural networks for power modeling for cloud and HPC: It’s worth it! ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5 (3):12:1–12:36, November 2020.

CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3388322.

Cassell:2018:DPM

[CSS^{+}18] Benjamin Cassell, Tyler Szepesi, Jim Summers, Tim Brecht, Derek Eager, and Bernard Wong. Disk prefetching mechanisms for increasing HTTP streaming video server throughput. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 3(2):7:1–7:??, April 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3164536.

Chen:2019:EQA

[CZCC19] Hao Chen, Yijia Zhang, Michael C. Caramanis, and Ayse K. Coskun. EnergyQARE: QoS-aware data center participation in smart grid regulation service reserve provision. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(1):2:1– 2:??, March 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3243172.

Chen:2018:CEO

[CZL^{+}18] Shutong Chen, Zhi Zhou, Fangming Liu, Zongpeng Li, and Shaolei Ren. CloudHeat: An efficient online market mechanism for datacenter heat harvesting. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(3):11:1–11:??, August 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3199675.

Du:2017:SCB

[DD17] Yuhuan Du and Gustavo De Veciana. Scheduling for cloudbased computing systems to support soft real-time applications. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(3):13:1–13:??, September 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3063713.

Du:2018:EOL

[DD18] Yuhuan Du and Gustavo De Veciana. Efficiency and optimality of largest deficit first prioritization: Dynamic user prioritization for soft real-time applications. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(3):10:1–10:??, August 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3200479.

DeCicco:2016:CTA

[DGRL16] Luca De Cicco, Yixi Gong, Dario Rossi, and Emilio Leonardi.

A control-theoretic analysis of low-priority congestion control reprioritization under AQM. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):17:1–17:33, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2934652.

Domingues:2021:RHC

[DMD^{+}21] Guilherme Domingues, Gabriel Mendonça, Edmundo De Souza E.Silva, Rosa M. M. Leão, Daniel S. Menasché, Ori Rottenstreich, Mostafa Dehghan, and Don Towsley. The role of hysteresis in caching systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 6(1):2:1–2:38, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450564.

Dong:2017:CAT

[DWS17] Fang Dong, Kui Wu, and Venkatesh Srinivasan. Copula analysis of temporal dependence structure in Markov modulated Poisson process and its applications. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(3):14:1–14:??, September 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3089254.

Engelmann:2021:CRA

[EJ21] Anna Engelmann and Admela Jukan. A combinatorial reliability analysis of generic service function chains in data center networks. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(3): 9:1–9:24, September 2021. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477046.

Friedlander:2019:GLC

[FA19] Eric Friedlander and Vaneet Aggarwal. Generalization of LRU cache replacement policy with applications to video streaming. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(3):18:1–18:??, September 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3345022.

Fu:2016:FPP

[FB16] Yongquan Fu and Ernst Biersack. False-positive probability and compression optimization for tree-structured Bloom filters. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):19:1–19:39, September 2016. CODEN ????

ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2940324.

Fan:2021:PAI

[FGK^{+}21] Caixiang Fan, Sara Ghaemi, Hamzeh Khazaei, Yuxiang Chen, and Petr Musilek. Performance analysis of the IOTA DAG-based distributed ledger. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6 (3):10:1–10:20, September 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3485188.

Fricker:2016:AOS

[FGRT16] Christine Fricker, Fabrice Guillemin, Philippe Robert, and Guilherme Thompson. Analysis of an offloading scheme for data centers in the framework of fog computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):16:1–16:18, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2950047.

Ferragut:2017:CVC

[FPW17] Andres Ferragut, Fernando Paganini, and Adam Wierman. Controlling the variability of capacity allocations using service deferrals. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(3):15:1–15:??, September 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3086506.

Gaeta:2021:MNI

[GG21] Rossano Gaeta and Marco Grangetto. Malicious node identification in coded distributed storage systems under pollution attacks. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(3):12:1– 12:27, September 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3491062.

Gupta:2020:SPD

[GHV20] Manu K. Gupta, N. Hemachandra, and J. Venkateswaran. Some parameterized dynamic priority policies for two-class M/G/1 queues: Completeness and applications. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(2):10:1– 10:37, April 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3384390.

Geissler:2021:DTM

[GLL^{+}21] Stefan Geissler, Stanislav Lange, Leonardo Linguaglossa, Dario Rossi, Thomas Zinner, and Tobias Hossfeld. Discrete-time modeling of NFV accelerators that exploit batched processing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(3):11:1–11:27, September 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3488243.

Garetto:2016:UAP

[GLM16] Michele Garetto, Emilio Leonardi, and Valentina Martina. A unified approach to the performance analysis of caching systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(3):12:1– 12:28, May 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2896380.

Gupta:2019:ERS

[GM19] Samarth Gupta and Sharayu Moharir. Effect of recommendations on serving content with unknown demand. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(1):4:1– 4:??, March 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3289324.

Golubchik:2021:MNE

[Gol21] Leana Golubchik. A message from the new Editor-in-Chief.

ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 5(4):15:1, March 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3432597.

Ghaderi:2016:SSS

[GSS16] Javad Ghaderi, Sanjay Shakkottai, and R. Srikant. Scheduling storms and streams in the cloud. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):14:1–14:28, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2904080.

Herbst:2018:QCP

[HBK^{+}18] Nikolas Herbst, André Bauer, Samuel Kounev, Giorgos Oikonomou, Erwin Van Eyk, George Kousiouris, Athanasia Evangelinou, Rouven Krebs, Tim Brecht, Cristina L. Abad, and Alexandru Iosup. Quantifying cloud performance and dependability: Taxonomy, metric design, and emerging challenges. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3 (4):19:1–19:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/citation.cfm?id=3236332.

Harrison:2016:EPT

[HPK16] Peter G. Harrison, Naresh M. Patel, and William J. Knottenbelt. Energy–performance tradeoffs via the EP queue. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1 (2):6:1–6:31, June 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

                                        http://dl.acm.org/citation.
                                        cfm?id=2818726.

Harrison:2019:MRT

[HPPQ19] P. G. Harrison, N. M. Patel, J. F. Pérez, and Z. Qiu. Managing response time tails by sharding. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(1):5:1– 5:??, March 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3300143.

Hellemans:2019:PRD

[HV19] Tim Hellemans and Benny Van Houdt. Performance of redundancy(d) with identical/ independent replicas. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4 (2):9:1–9:??, June 2019. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/citation.cfm?id=3316768.

Izadpanah:2019:PAP

[IADB19] Ramin Izadpanah, Benjamin A.

Allan, Damian Dechev, and Jim Brandt. Production application performance data streaming for system monitoring. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(2):8:1– 8:??, June 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3319498.

Ilyushkin:2018:EPE

[IAEH^{+}18] Alexey Ilyushkin, Ahmed Ali-Eldin, Nikolas Herbst, André Bauer, Alessandro V. Papadopoulos, Dick Epema, and Alexandru Iosup. An experimental performance evaluation of autoscalers for complex workflows. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(2):8:1– 8:??, April 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3164537.

Izagirre:2016:STA

[IAV16] A. Izagirre, U. Ayesta, and I. M. Verloop. Sojourn time approximations for a discriminatory processor sharing queue. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 1(1):5:1–5:31, March 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2812807.

Jiang:2020:LHS

[JLZ20] Qingye Jiang, Young Choon Lee, and Albert Y. Zomaya. The limit of horizontal scaling in public clouds. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(1):6:1– 6:22, February 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

                                        https://dl.acm.org/doi/abs/
                                        10.1145/3373356.

Jiang:2018:CTA

[JNT18] Bo Jiang, Philippe Nain, and Don Towsley. On the convergence of the TTL approximation for an LRU cache under independent stationary request processes. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(4):20:1–20:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3239164.

Jiang:2021:CCS

[JNT21] Bo Jiang, Philippe Nain, and Don Towsley. Covert cycle stealing in a single FIFO server. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 6(2):5:1–5:33, June 2021.

CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3462774.

Joshi:2017:ERT

[JSW17] Gauri Joshi, Emina Soljanin, and Gregory Wornell. Efficient redundancy techniques for latency reduction in cloud systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(2):12:1– 12:30, May 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
3055281.

Khan:2018:RAE

[KHN^{+}18] Kashif Nizam Khan, Mikael Hirki, Tapio Niemi, Jukka K. Nurminen, and Zhonghong Ou. RAPL in action: Experiences in using RAPL for power measurements. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(2):9:1– 9:??, April 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3177754.

Kalbasi:2019:AAM

[KKR19] Amir Kalbasi, Diwakar Krishnamurthy, and Jerry Rolia. AMIR: Analytic method for improving responsiveness by reducing burstiness. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(4):19:1– 19:36, December 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/abs/
10.1145/3365669.

KhudaBukhsh:2019:PPE

[KKRK19] Wasiur R. KhudaBukhsh, Sounak Kar, Amr Rizk, and Heinz Koeppl. Provisioning and performance evaluation of parallel systems with output synchronization. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(1):6:1– 6:??, March 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3300142.

Koziolek:2018:SIS

[KS18] Anne Koziolek and Evgenia Smirni. Special issue: Selected paper from the 8th ACM/ SPEC International Conference on Performance Engineering (ICPE 2017). ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(2):6:1– 6:??, April 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3186329.

Kelley:2017:OMA

[KSM^{+}17] Jaimie Kelley, Christopher Stewart, Nathaniel Morris, Devesh Tiwari, Yuxiong He, and Sameh Elnikety. Obtaining and managing answer quality for online data-intensive services. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(2):11:1–11:31, May 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
3055280.

Krishnasamy:2016:DSR

[KSSO16] Subhashini Krishnasamy, Rajat Sen, Sanjay Shakkottai, and Sewoong Oh. Detecting sponsored recommendations. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(1):6:1– 6:29, November 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2988543.

Lashgar:2016:ESM

[LB16] Ahmad Lashgar and Amirali Baniasadi. Employing softwaremanaged caches in OpenACC: Opportunities and benefits. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 1(1):2:1–2:34, March 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2798724.

Liu:2017:FPE

[LCD^{+}17] Yanpei Liu, Guilherme Cox, Qingyuan Deng, Stark C. Draper, and Ricardo Bianchini. Fast power and energy management for future many-core systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(3):17:1–17:??, September 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3086504.

Liu:2021:DES

[LKC^{+}21] Zhengchun Liu, Rajkumar Kettimuthu, Joaquin Chung, Rachana Ananthakrishnan, Michael Link, and Ian Foster. Design and evaluation of a simple data interface for efficient data transfer across diverse storage. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(1): 3:1–3:25, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3452007.

Liu:2018:BGB

[LLTL18] Chubo Liu, Kenli Li, Zhuo Tang, and Keqin Li. Bargaining gamebased scheduling for performance guarantees in cloud computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 3(1):1:1–1:??, February 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3141233.

Lim:2019:PCI

[LLW^{+}19] Chiun Lin Lim, Ki Suh Lee, Han Wang, Hakim Weatherspoon, and Ao Tang. Packet clustering introduced by routers: Modeling, analysis, and experiments. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(3):15:1–15:??, September 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3345032.

Li:2018:MFA

[LRS18] Bin Li, Aditya Ramamoorthy, and R. Srikant. Mean-field analysis of coding versus replication in large data storage systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 3(1):3:1–3:??, February 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3159172.

Liu:2017:ECE

[LS17] Yu-Hang Liu and Xian-He Sun.

Evaluating the combined effect of memory capacity and concurrency for many-core chip design. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(2):9:1– 9:25, May 2017. CODEN ????

ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        3038915.

Liao:2020:TEB

[LSC^{+}20] Jianwei Liao, Zhibing Sha, Zhigang Cai, Zhiming Liu, Kenli Li, Wei-Keng Liao, Alok N. Choudhary, and Yutaka Ishiakwa. Toward efficient block replication management in distributed storage. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5 (3):13:1–13:27, November 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3412450.

Li:2018:MFG

[LXG^{+}18] Jian Li, Bainan Xia, Xinbo Geng, Hao Ming, Srinivas Shakkottai, Vijay Subramanian, and Le Xie. Mean field games in nudge systems for societal networks. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(4):15:1–15:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3232076.

Li:2017:IPO

[LXW^{+}17] Hao Li, Xinhai Xu, Miao Wang, Chao Li, Xiaoguang Ren, and Xuejun Yang. Insertion of PETSc in the Open-FOAM framework. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2 (3):16:1–16:??, September 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/citation.cfm?id=3098821.

Li:2019:QCS

[LZL^{+}19] Zhenhua Li, Yongfeng Zhang, Yunhao Liu, Tianyin Xu, Ennan Zhai, Yao Liu, Xiaobo Ma, and Zhenyu Li. A quantitative and comparative study of network-level efficiency for cloud storage services. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(1):3:1– 3:??, March 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3274526.

Liu:2016:SSS

[LZW^{+}16] Qingyun Liu, Xiaohan Zhao, Walter Willinger, Xiao Wang, Ben Y. Zhao, and Haitao Zheng. Self-similarity in social network dynamics. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(1):5:1– 5:26, November 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2994142.

Molka:2016:CAW

[MC16] Karsten Molka and Giuliano Casale. Contention-aware workload placement for in-memory databases in cloud environments. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(1):1:1–1:29, November 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2961888.

Mciver:2020:ISS

[MH20] Annabelle Mciver and András Horváth. Introduction to the special section on quantitative evaluation of systems (QEST 2018). ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(1):1:1, February 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.1145/
                                        3376999.

Marques:2021:MRM

[MIS21] Diogo Marques, Aleksandar Ilic, and Leonel Sousa. Mansard roofline model: Reinforcing the accuracy of the roofs. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6 (2):7:1–7:23, June 2021. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3475866.

Malik:2018:SAL

[MRH18] Maria Malik, Setareh Rafatirad, and Houman Homayoun. System and architecture level characterization of big data applications on big and little core server architectures. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(3):14:1–14:??, August 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3229049.

Marin:2020:DRA

[MRS20] Andrea Marin, Sabina Rossi, and Matteo Sottana. Dynamic resource allocation in fork–join queues. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(1):3:1– 3:28, February 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/abs/
10.1145/3372376.

Makrani:2021:APM

[MSN^{+}21] Hosein Mohamamdi Makrani, Hossein Sayadi, Najmeh Nazari, Sai Mnoj Pudukotai Dinakarrao, Avesta Sasan, Tinoosh Mohsenin, Setareh Rafatirad, and Houman Homayoun. Adaptive performance modeling of data-intensive workloads for resource provisioning in virtualized environment. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(4): 18:1–18:24, March 2021. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3442696.

Nemati:2021:HBV

[NASD21] Hani Nemati, Seyed Vahid Azhari, Mahsa Shakeri, and Michel Dagenais. Host-based virtual machine workload characterization using hypervisor trace mining. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(1): 4:1–4:25, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

                                        https://dl.acm.org/doi/10.
                                        1145/3460197.

Nguyen:2019:PFR

[NBP19] T. T. Hang Nguyen, Olivier Brun, and Balakrishna J. Prabhu. Performance of a fixed reward incentive scheme for two-hop DTNs with competing relays. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(2):12:1– 12:??, June 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3325288.

Neglia:2017:ATA

[NCF^{+}17] Giovanni Neglia, Damiano Carra, Mingdong Feng, Vaishnav Janardhan, Pietro Michiardi, and Dimitra Tsigkari. Accesstime-aware cache algorithms. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(4):21:1–21:??, December 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3149001.

Narayana:2021:RER

[NMK21] V. S. Ch Lakshmi Narayana, Sharayu Moharir, and Nikhil Karamchandani. On renting edge resources for service hosting. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(2):8:1– 8:30, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3478433.

Namazi:2019:SSO

[NSMA19] Alireza Namazi, Saeed Safari, Siamak Mohammadi, and Meisam Abdollahi. SORT: Semi online reliable task mapping for embedded multi-core systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 4(2):11:1–11:??, June 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3322899.

Nain:2016:FDD

[NT16] Philippe Nain and Don Towsley.

File dissemination in dynamic graphs: The case of independent and correlated links in series.

ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(1):4:1–4:23, November 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2981344.

Nordio:2018:STQ

[NTLM18] Alessandro Nordio, Alberto Tarable, Emilio Leonardi, and Marco Ajmone Marsan. Selecting the top-quality item through crowd scoring. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(1):2:1–2:??, February 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

                                        tronic). URL https://dl.acm.
                                        org/citation.cfm?id=3157736.

Nasiriani:2016:FAC

[NWK^{+}16] Neda Nasiriani, Cheng Wang, George Kesidis, Bhuvan Urgaonkar, Lydia Y. Chen, and Robert Birke. On fair attribution of costs under peak-based pricing to cloud tenants. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(1):3:1– 3:28, November 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

                                        http://dl.acm.org/citation.
                                        cfm?id=2970815.

Niu:2017:RAS

[NXL17] Di Niu, Hong Xu, and Baochun Li. Resource auto-scaling and sparse content replication for video storage systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(4):19:1–19:??, December 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3079045.

Papadopoulos:2016:PPE

[PAEÅ^{+}16] Alessandro Vittorio Papadopoulos, Ahmed Ali-Eldin, Karl-Erik Årzén, Johan Tordsson, and Erik Elmroth. PEAS: A performance evaluation framework for autoscaling strategies in cloud applications. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):15:1–15:31, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2930659.

Pajevic:2018:EPC

[PFK18] Ljubica Pajevic, Viktoria Fodor, and Gunnar Karlsson. Ensuring persistent content in opportunistic networks via stochastic stability analysis. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3 (4):16:1–16:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/citation.cfm?id=3232161.

Phan:2019:NFE

[PPIR19] Tien-Dat Phan, Guillaume Pallez, Shadi Ibrahim, and Padma Raghavan. A new framework for evaluating straggler detection mechanisms in MapReduce. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(3):14:1–14:??, September 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3328740.

Petsas:2017:MMA

[PPP^{+}17] Thanasis Petsas, Antonis Papadogiannakis, Michalis Polychronakis, Evangelos P. Markatos, and Thomas Karagiannis. Measurement, modeling, and analysis of the mobile app ecosystem. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(2):7:1–7:33, May 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

                                        dl.acm.org/citation.cfm?id=
                                        2993419.

Plakia:2020:SSS

[PTA^{+}20] Maria Plakia, Evripides Tzamousis, Thomais Asvestopoulou, Giorgos Pantermakis, Nick Filippakis, Henning Schulzrinne, Yana Kane-Esrig, and Maria Papadopouli. Should I stay or should I go: Analysis of the impact of application QoS on user engagement in YouTube. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5 (2):9:1–9:32, April 2020. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377873.

Qin:2021:OOA

[QE21] Tiancheng Qin and S. Rasoul Etesami. Optimal online algorithms for file-bundle caching and generalization to distributed caching. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6(1):1:1– 1:23, March 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3445028.

Rattaro:2020:QPD

[RAMB20] Claudina Rattaro, Laura Aspirot, Ernesto Mordecki, and Pablo Belzarena. QoS provision in a dynamic channel allocation based on admission control decisions. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(1):5:1– 5:29, February 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/abs/
10.1145/3372786.

Raeis:2020:AQM

[RBL20] Majid Raeis, Almut Burchard, and Jörg Liebeherr. Analysis of a queueing model for energy storage systems with selfdischarge. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(3):14:1– 14:26, November 2020. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3422711.

Roos:2016:DDE

[RS16] Stefanie Roos and Thorsten Strufe. Dealing with dead ends: Efficient routing in darknets. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 1(1):4:1–4:30, March 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2809779.

Ray:2018:SSC

[RSS18] Avik Ray, Sujay Sanghavi, and Sanjay Shakkottai. Searching for a single community in a graph. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(3):13:1–13:??, August 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3200863.

Sonenberg:2021:PAW

[SKV21] Nikki Sonenberg, Grzegorz Kielanski, and Benny Van Houdt. Performance analysis of work stealing in large-scale multithreaded computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 6 (2):6:1–6:28, June 2021. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3470887.

Skevakis:2019:SOF

[SLH19] Emmanouil Skevakis, Ioannis Lambadaris, and Hassan Halabian. Scheduling for optimal file-transfer delay using chunked random linear network coding broadcast. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 4(3):17:1–17:??, September 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3340242.

Simhon:2017:ARG

[SS17] Eran Simhon and David Starobinski. Advance reservation games. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 2(2):10:1–10:21, May 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
3053046.

Sieber:2020:SAU

[SSB^{+}20] Christian Sieber, Susanna Schwarzmann, Andreas Blenk, Thomas Zinner, and Wolfgang Kellerer. Scalable application- and useraware resource allocation in enterprise networks using endhost pacing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(3):11:1– 11:41, November 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/10.
1145/3381996.

Salamati:2020:LAT

[SSM20] Mahmoud Salamati, Sadegh Soudjani, and Rupak Majumdar. A Lyapunov approach for timebounded reachability of CTMCs and CTMDPs. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(1):2:1– 2:29, February 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/abs/
10.1145/3371923.

Tavakkol:2016:PED

[TMASA16] Arash Tavakkol, Pooyan Mehrvarzy, Mohammad Arjomand, and Hamid Sarbazi-Azad. Performance evaluation of dynamic page allocation strategies in SSDs. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(2):7:1– 7:33, June 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2829974.

Towsley:2016:I

[TW16] Don Towsley and Carey Williamson.

Introduction. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(1): 1:1, March 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2893179.

Varki:2018:GGP

[Var18] Elizabeth Varki. GPSonflow:

Geographic positioning of storage for optimal nice flow. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 3(3):12:1–12:??, August 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3197656.

Vassio:2020:UIO

[VGCL20] Luca Vassio, Michele Garetto, Carla Chiasserini, and Emilio Leonardi. User interaction with online advertisements: Temporal modeling and optimization of ads placement. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5 (2):8:1–8:26, April 2020. CO-DEN ???? ISSN 2376-3639 (print), 2376-3647 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3377144.

Vergara:2016:FIC

[VNTA16] Ekhiotz Jon Vergara, Simin Nadjm-Tehrani, and Mikael Asplund. Fairness and incentive considerations in energy apportionment policies. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(1):2:1– 2:29, November 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2970816.

Vial:2019:RCP

[VS19] Daniel Vial and Vijay Subramanian. On the role of clustering in personalized PageRank estimation. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(4):21:1–21:33, December 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/
3366635.

Verschoren:2019:EDC

[VV19] Robin Verschoren and Benny Van Houdt. On the endurance of the d-choices garbage collection algorithm for flash-based SSDs. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(3):13:1–13:??, September 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3326121.

Wang:2018:QMS

[WCKN18] Weikun Wang, Giuliano Casale, Ajay Kattepur, and Manoj K. Nambiar. QMLE: A methodology for statistical inference of service demands from queueing data. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(4):17:1–17:??, September 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3233180.

Wu:2019:FAS

[WDGC19] Xiaohu Wu, Francesco De Pellegrini, Guanyu Gao, and Giuliano Casale. A framework for allocating server time to spot and ondemand services in cloud computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(4):20:1–20:31, December 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/doi/abs/10.1145/
3366682.

Wang:2019:ESR

[WJW19] Da Wang, Gauri Joshi, and Gregory W. Wornell. Efficient straggler replication in large-scale parallel computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(2):7:1– 7:??, June 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3310336.

Wang:2018:EAA

[WLU18] Cheng Wang, Qianlin Liang, and Bhuvan Urgaonkar. An empirical analysis of Amazon EC2 spot instance features affecting cost-effective resource procurement. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 3(2):6:1– 6:??, April 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/citation.cfm?id=
                                        3164538.

Wang:2019:CSC

[WXL^{+}19] Xiaoming Wang, Di Xiao, Xiaoyong Li, Daren B. H. Cline, and Dmitri Loguinov. Consistent sampling of churn under periodic non-stationary arrivals in distributed systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 4(4):22:1–22:33, December 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

                                        dl.acm.org/doi/abs/10.1145/
                                        3368510.

Wang:2019:OTA

[WZK^{+}19] Qingyang Wang, Shungeng Zhang, Yasuhiko Kanemasa, Calton Pu, Balaji Palanisamy, Lilian Harada, and Motoyuki Kawaba. Optimizing N-tier application scalability in the cloud: A study of soft resource allocation. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(2):10:1– 10:??, June 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3326120.

Xie:2016:DAI

[XLT16] Hong Xie, John C. S. Lui, and Don Towsley. Design and analysis of incentive and reputation mechanisms for online crowdsourcing systems. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(3): 13:1–13:27, May 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

http://dl.acm.org/citation.
cfm?id=2897510.

Yang:2020:EPC

[YLL20] Jinfeng Yang, Bingzhe Li, and David J. Lilja. Exploring performance characteristics of the Optane 3D Xpoint storage technology. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 5(1):4:1– 4:28, February 2020. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL

https://dl.acm.org/doi/abs/
10.1145/3372783.

Yi:2017:CDC

[YLN^{+}17] Xiaomeng Yi, Fangming Liu, Di Niu, Hai Jin, and John C. S.

Lui. Cocoa: Dynamic containerbased group buying strategies for cloud computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 2(2):8:1– 8:31, May 2017. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
3022876.

Yan:2016:PAF

[YMRS16] Feng Yan, Xenia Mountrouidou, Alma Riska, and Evgenia Smirni. PREFiguRE: An analytic framework for HDD management. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 1(3):10:1–10:27, May 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2872331.

Yin:2019:ETL

[YYX^{+}19] Ping Yin, Sen Yang, Jun Xu, Jim Dai, and Bill Lin. Efficient traffic load-balancing via incremental expansion of routing choices. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 4(1):1:1– 1:??, March 2019. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3243173.

Zhou:2018:OED

[ZLW18] Ruiting Zhou, Zongpeng Li, and Chuan Wu. An online emergency demand response mechanism for cloud computing. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOM-PECS), 3(1):5:1–5:??, February 2018. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL https://

dl.acm.org/citation.cfm?id=
3177755.

Zhang:2016:TIM

[ZRWL16] Linquan Zhang, Shaolei Ren, Chuan Wu, and Zongpeng Li. A truthful incentive mechanism for emergency demand response in geo-distributed colocation data centers. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(4):18:1–18:23, September 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2950046.

Zhang:2016:VSL

[ZWHD16] Deli Zhang, Jeremiah Wilke, Gilbert Hendry, and Damian Dechev. Validating the simulation of large-scale parallel applications using statistical characteristics. ACM Transactions on Modeling and Performance Evaluation of Computing Systems (TOMPECS), 1(1):3:1– 3:22, March 2016. CODEN ???? ISSN 2376-3639 (print), 2376-3647 (electronic). URL http://

dl.acm.org/citation.cfm?id=
2809778.